# Motivic classes of stacks in finite characteristic and applications to stacks of Higgs bundles

### v1

@@ 2.2. Dimensional completion
Following Behrend and Dhillon in [BD07, Section 2.1], we define the dimensional completion of $\operatorname{Mot}(k)$.
@@
This is used in [FSS18, Section 4] to prove the “motivic Harder formula”.
@@
2.3. Motivic zeta function and pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$
The main goal of this section is to define the pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$ and the main result is Theorem 2.18. We recall the definitions of quotient by finite groups following [Mus11, Section A.1] and motivic zeta function from [Kap00, (1.3)]. Then we recall the definition of pre-$\lambda$ structure, see [Knu73, Chapter I]. In this paper we only consider pre-$\lambda$ structures on commutative rings with units.
@@ 2.3. Motivic zeta function and pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$
Let $X$ be a scheme of finite type over a field $k$ and let $G$ be an abstract finite group, acting on $X$ by algebraic automorphisms over $k$ (we make the convention that when $X=\operatorname{Spec}A$, the group $G$ acts on the ring $A$ on the left as $g\cdot a$, so it acts on $\operatorname{Spec}A$ on the right respectively). We denote by $\sigma_{g}$ the automorphism of $X$ corresponding to $g\in G$. A quotient of $X$ by $G$ is a morphism $\pi\colon X\to W$ with the following two properties:
@@
    $\pi$ is $G$-invariant, that is $\pi\circ\sigma_{g}=\pi$ for every $g\in G$;
@@
See [Mus11, Section A.1], especially the discussions after [Mus11, Corollary A.3]. ∎
@@
Let $X$ be a quasi-projective scheme over $k$, and $G$ an abstract finite group acting on $X$ by algebraic automorphisms over $k$. Let $\pi\colon X\to X/G$ be the quotient morphism. Let $X=\bigsqcup_{i}X_{i}$ be a stratification with $G$-invariant locally closed subschemes $X_{i}$. We have in $\operatorname{Mot}_{sch}(k)$
@@
Let $X=\bigsqcup_{i}X_{i}$ be a stratification of $X$ with $G$-invariant locally closed subschemes $X_{i}$. We have
@@
We define the motivic zeta function for every reduced quasi-projective scheme $X$:
$\zeta_{X}(z)=\sum_{n\geq 0}[Sym^{n}X]\cdot z^{n}\in 1+z\cdot\operatorname{Mot}_{sch}(k)[[z]],$
@@
For $k=\mathbb{F}_{q}$ a finite field with $q$ elements, the image of the series $\zeta_{X}(z)$ under the counting measure in Proposition 3.2 coincides with the local zeta function, see [Mus11, Proposition 7.31].
@@
    $\zeta_{X}(z)=\zeta_{Y}(z)\zeta_{X-Y}(z)$;
@@
    $\zeta_{X\times\mathbb{A}_{k}^{n}}(z)=\zeta_{X}(\mathbb{L}^{n}z)$.
@@
The two statements are well-known (see, e.g., [Mus11, Propositions 7.28 and 7.32]). To illustrate the importance of the second relation of Definition 2.1, we sketch a proof of (i) following [Mus11, Proposition 7.28 and Lemma 7.29]). If $k$ has characteristic 0, the proof of (ii) is also given in [GZLMH04, Statement 3].
For fixed $n$, let $i,j$ be nonnegative integers such that $i+j=n$. Denote by $W^{i,j}$ the locally closed subset of $X^{n}$ given by $\cup_{g\in S_{n}}(Y^{i}\times(X-Y)^{j})g$.
@@
We recall the definition of pre-$\lambda$ structure, see [Knu73, Chapter I].
@@
is called a pre-$\lambda$ structure. In other words, for all $a,b\in R$ we have
@@
    $\lambda_{a}(z)\equiv 1+az\ (\text{mod}\ z^{2})$;
@@
Now we want to define a pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$ similarly to [FSS18, Section 1.3]. First, we need a lemma.
@@
Given a pre-$\lambda$ structure on a ring $R$ and $l\in R\backslash\{0\}$, if for all $r\in R$ we have
@@
then we have a unique pre-$\lambda$ structure on $R[l^{-1}]$ such that the homomorphism $R\to R[l^{-1}]$ is a homomorphism of pre-$\lambda$-rings.
@@
It is clear that we obtained a pre-$\lambda$ structure on $R[l^{-1}]$ and the homomorphism $R\to R[l^{-1}]$ is a homomorphism of pre-$\lambda$-rings. ∎
For any field $k$, we define $\operatorname{Mot}^{qpr}_{sch}(k)$ as the abelian group generated by isomorphism classes of reduced quasi-projective schemes of finite type over $k$ modulo relations of Definition 2.3. We have a group homomorphism $\Phi\colon\operatorname{Mot}^{qpr}_{sch}(k)\to\operatorname{Mot}_{sch}(k)$ such that $\Phi([X])=[X]$ for any reduced quasiprojective scheme $X$. Indeed this is an isomorphism.
@@
There is a unique continuous pre-$\lambda$ structure on $\overline{\operatorname{Mot}}(k)$ that sends the class of a reduced quasi-projective scheme $X$ to $(\zeta_{X}(-z))^{-1}$ and for any $A\in\overline{\operatorname{Mot}}(k)$ and any $n\in\mathbb{Z}$ we have $\zeta_{\mathbb{L}^{n}A}(z)=\zeta_{A}(\mathbb{L}^{n}z)$.
@@
Similarly to [Mus11, Proposition 7.28], a pre-$\lambda$ structure can be defined on $\operatorname{Mot}^{qpr}_{sch}(k)$ by
@@
By Lemma 2.17, we first obtain a pre-$\lambda$ structure on $\operatorname{Mot}_{sch}(k)$. We extend the pre-$\lambda$ structure on $\operatorname{Mot}_{sch}(k)$ to the pre-$\lambda$ structure on $\operatorname{Mot}_{sch}(k)[\mathbb{L}^{-1}]$ by Proposition 2.14(ii) and Lemma 2.16. Next we extend the pre-$\lambda$ structure to $\overline{\operatorname{Mot}}_{sch}(k)$ by continuity. Finally, by Proposition 2.9, we have an isomorphism $\overline{\operatorname{Mot}}(k)\simeq\overline{\operatorname{Mot}}_{sch}(k).$
@@
$\lambda_{A}(z)=\lim_{i\to\infty}\frac{\zeta_{Z_{i}}(-\mathbb{L}^{n_{i}}z)}{\zeta_{Y_{i}}(-\mathbb{L}^{n_{i}}z)},$
@@ 2.4. Stack motivic functions
We define first the group of motivic functions for stacks of finite type similarly to the definition in [FSS18, Section 2] except that we add a relation for the surjective and universally injective morphisms.
@@
Let $\mathcal{X}$ be an Artin stack of finite type. The abelian group $\operatorname{Mot}(\mathcal{X})$ is the group generated by the isomorphism classes of finite type morphisms $\mathcal{Y}\to\mathcal{X}$ modulo relations
@@
    $\mathcal{Y}_{1}\to\mathcal{X}]=[\mathcal{Y}_{2}\to\mathcal{X}]+[(\mathcal{Y}_{1}-\mathcal{Y}_{2})\to\mathcal{X}]$ whenever $\mathcal{Y}_{2}$ is a closed substack of $\mathcal{Y}_{1}$;
@@
    $\mathcal{Y}_{2}\to\mathcal{X}]=[\mathcal{Y}_{1}\to\mathcal{X}]$ if there is a surjective and universally injective morphism $\mathcal{Y}_{2}\to\mathcal{Y}_{1}$ of stacks over $\mathcal{X}$;
@@
    It $\pi\colon\mathcal{Y}_{1}\to\mathcal{X}$ is a morphism of finite type and $\psi\colon\mathcal{Y}_{2}\to\mathcal{Y}_{1}$ is a vector bundle of rank $r$, then
    $\mathcal{Y}_{2}\xrightarrow{\pi\circ\psi}\mathcal{X}]=[\mathcal{Y}_{1}\times\mathbb{A}^{r}_{k}\xrightarrow{\pi\circ p_{1}}\mathcal{X}].$
@@
We define the pullback and pushforward homomorphisms for stacks of finite type. By [SP18, Tag03MH and Tag03MW], the surjective and universally injective morphisms are stable under base change. Similarly to the content of [FSS18, Section 2.1], for a morphism $f\colon\mathcal{X}\to\mathcal{Y}$ of stacks of finite type possibly over different fields, we define the pullback homomorphism $f^{\ast}\colon\operatorname{Mot}(\mathcal{Y})\to\operatorname{Mot}(\mathcal{X})$ as
@@
We also have the pushforward homomorphism $f_{!}\colon\operatorname{Mot}(\mathcal{X})\to\operatorname{Mot}(\mathcal{Y})$ such that
@@
The pullbacks and pushforwards satisfy the same properties as in [FSS18, Section 2.1]. For $k$-stacks $\mathcal{X}$ and $\mathcal{Y}$ of finite type, we have an external product $\boxtimes\colon\operatorname{Mot}(\mathcal{X})\otimes_{\mathbb{Z}}\operatorname{Mot}(\mathcal{Y})\to\operatorname{Mot}(\mathcal{X}\times\mathcal{Y})$.
@@
We define
@@
(2.1) $\operatorname{Mot}^{fin}(\mathcal{X})\coloneqq\cup_{\mathcal{U}\in\operatorname{Opf}(\mathcal{X})}(j_{\mathcal{U}})_{!}\operatorname{Mot}(\mathcal{U})\subset\operatorname{Mot}(\mathcal{X}),$
@@
where the limit is taken over the partially ordered set $\operatorname{Opf}(\mathcal{X})$. Note that the pullback $f^{\ast}$ and pushforward $f_{!}$ can be extended to the completions. The group $\overline{\operatorname{Mot}}^{fin}(\mathcal{X})$ can also be defined similarly to (2.1).

### v2

@@ 2.2. Dimensional completion
Following Behrend and Dhillon’s results in [BD07, Section 2.1], we define the dimensional completion of $\operatorname{Mot}(k)$.
@@
We note that calculations of motivic classes of various stacks associated with a curve are based on the “motivic Harder formula” proved in [FSS18, Section 4], this formula is only valid in the completed ring of motivic classes.
@@
2.3. Motivic zeta-function and pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$
The main goal of this section is to define the pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$ and the main result is Theorem 2.18. We recall the definitions of quotient by finite groups following [Mus11, Section A.1] and motivic zeta-function from [Kap00, (1.3)]. Then we recall the definition of pre-$\lambda$-ring structure, see [Knu73, Chapter I]. In this paper we only consider pre-$\lambda$-ring structures on commutative rings with units.
@@ 2.3. Motivic zeta-function and pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$
Let $X$ be a scheme of finite type over a field $k$ and let $G$ be an abstract finite group, acting on $X$ by algebraic automorphisms over $k$ (we make the convention that when $X=\operatorname{Spec}A$, the group $G$ acts on the ring $A$ on the left as $g\cdot a$, so it acts on $\operatorname{Spec}A$ on the right respectively). We denote by $\sigma_{g}$ the automorphism of $X$ corresponding to $g\in G$.
@@
A quotient of $X$ by $G$ is a morphism $\pi\colon X\to W$ with the following two properties:
@@
    $\pi$ is $G$-invariant, that is $\pi\circ\sigma_{g}=\pi$ for every $g\in G$.
@@
See [Mus11, Section A.1], especially the discussions of affine open covers after [Mus11, Corollary A.3]. ∎
@@
Let $X$ be a quasi-projective scheme over $k$, and $G$ an abstract finite group acting on $X$ by algebraic automorphisms over $k$. Let $X=\bigsqcup_{i}X_{i}$ be a stratification with $G$-invariant locally closed subschemes $X_{i}$. We have in $\operatorname{Mot}_{sch}(k)$
@@
Let $X=\bigsqcup_{i}X_{i}$ be a stratification of $X$ with $G$-invariant locally closed subschemes $X_{i}$. Let $\pi\colon X\to X/G$ be the quotient morphism. We have
@@
We define the motivic zeta-function for every reduced quasi-projective scheme $X$:
$\zeta_{X}(z)\coloneqq\sum_{n\geq 0}[Sym^{n}X]\cdot z^{n}\in 1+z\cdot\operatorname{Mot}_{sch}(k)[[z]],$
@@
For $k=\mathbb{F}_{q}$ a finite field with $q$ elements, the image of the series $\zeta_{X}(z)$ under the counting measure in Theorem 3.2 coincides with the local zeta-function, see [Mus11, Proposition 7.31].
@@
    $\zeta_{X}(z)=\zeta_{Y}(z)\zeta_{X-Y}(z)$.
@@
    $\zeta_{X\times_{k}\mathbb{A}_{k}^{n}}(z)=\zeta_{X}(\mathbb{L}^{n}z)$.
@@
This material is well-known (see e.g., [Mus11, Propositions 7.28 and 7.32]). To illustrate the importance of the second relation of Definition 2.1, we sketch a proof of (i) following [Mus11, Proposition 7.28 and Lemma 7.29]). If $k$ has characteristic 0, the proof of (ii) is also given in [GZLMH04, Statement 3].
For fixed $n$, let $i,j$ be nonnegative integers such that $i+j=n$. Denote by $W^{i,j}$ the locally closed subset of $X^{n}$ given by $\cup_{g\in S_{n}}(Y^{i}\times_{k}(X-Y)^{j})g$.
@@
We recall the definition of pre-$\lambda$-ring structure, see [Knu73, Chapter I].
@@
is called a pre-$\lambda$-ring structure. In other words, for all $a,b\in R$ we have
@@
    $\lambda_{a}(z)\equiv 1+az\ (\text{mod}\ z^{2})$.
@@
Now we want to define a pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$ similarly to [FSS18, Section 1.3]. First, we need a lemma.
@@
Given a pre-$\lambda$-ring structure on a ring $R$ and $l\in R\backslash\{0\}$, if for all $r\in R$ we have
@@
then we have a unique pre-$\lambda$-ring structure on $R[l^{-1}]$ such that the homomorphism $R\to R[l^{-1}]$ is a homomorphism of pre-$\lambda$-rings, i.e, the homomorphism commutes with the pre-$\lambda$-ring structure.
@@
It is clear that we obtained a pre-$\lambda$-ring structure on $R[l^{-1}]$ and the homomorphism $R\to R[l^{-1}]$ is a homomorphism of pre-$\lambda$-rings. ∎
For any field $k$, we define $\operatorname{Mot}^{qpr}_{sch}(k)$ as the abelian group generated by isomorphism classes of reduced quasi-projective schemes of finite type over $k$ modulo relations of Definition 2.3. We have a group homomorphism $\Phi\colon\operatorname{Mot}^{qpr}_{sch}(k)\to\operatorname{Mot}_{sch}(k)$ such that $\Phi([X])=[X]$ for any reduced quasi-projective scheme $X$. In fact, this is an isomorphism.
@@
There is a unique continuous pre-$\lambda$-ring structure on $\overline{\operatorname{Mot}}(k)$ that sends the class of a reduced quasi-projective scheme $X$ to $(\zeta_{X}(-z))^{-1}$ and for any $A\in\overline{\operatorname{Mot}}(k)$ and any $n\in\mathbb{Z}$ we have $\zeta_{\mathbb{L}^{n}A}(z)=\zeta_{A}(\mathbb{L}^{n}z)$.
@@
Similarly to [Mus11, Proposition 7.28], a pre-$\lambda$-ring structure can be defined on $\operatorname{Mot}^{qpr}_{sch}(k)$ by
@@
By Lemma 2.17, we first obtain a pre-$\lambda$-ring structure on $\operatorname{Mot}_{sch}(k)$. We extend the pre-$\lambda$-ring structure on $\operatorname{Mot}_{sch}(k)$ to the pre-$\lambda$-ring structure on $\operatorname{Mot}_{sch}(k)[\mathbb{L}^{-1}]$ by Proposition 2.14(ii) and Lemma 2.16. Next we extend the pre-$\lambda$-ring structure to $\overline{\operatorname{Mot}}_{sch}(k)$ by continuity. Finally, by Proposition 2.9, we have an isomorphism $\overline{\operatorname{Mot}}(k)\simeq\overline{\operatorname{Mot}}_{sch}(k).$
@@
$\lambda_{A}(z)=\lim_{i\to\infty}\frac{\zeta_{Z_{i}}(-\mathbb{L}^{-n_{i}}z)}{\zeta_{Y_{i}}(-\mathbb{L}^{-n_{i}}z)},$
@@
Remark 2.19.
We are using the opposite pre-$\lambda$-ring structure here because for the pre-$\lambda$-ring structure in [FSS18, Section 1.3], we have $\lambda(1)=(1-z)^{-1}$. Thus it admits no non-trivial homomorphism to a $\lambda$-ring (see Section 2.6).
@@ 2.4. Stack motivic functions
Following the definitions in [FSS18, Section 2], we define first the group of motivic functions for stacks of finite type except that we add a relation for the surjective and universally injective morphisms.
@@
Let $\mathcal{X}$ be an Artin stack of finite type over $k$. The abelian group $\operatorname{Mot}(\mathcal{X})$ is the group generated by the isomorphism classes of finite type morphisms $\mathcal{Y}\to\mathcal{X}$ modulo relations
@@
    $\mathcal{Y}_{1}\to\mathcal{X}]=[\mathcal{Y}_{2}\to\mathcal{X}]+[(\mathcal{Y}_{1}-\mathcal{Y}_{2})\to\mathcal{X}]$ whenever $\mathcal{Y}_{2}$ is a closed substack of $\mathcal{Y}_{1}$.
@@
    $\mathcal{Y}_{2}\to\mathcal{X}]=[\mathcal{Y}_{1}\to\mathcal{X}]$ if there is a surjective and universally injective morphism $\mathcal{Y}_{2}\to\mathcal{Y}_{1}$ of stacks over $\mathcal{X}$.
@@
    If $\pi\colon\mathcal{Y}_{1}\to\mathcal{X}$ is a morphism of finite type and $\psi\colon\mathcal{Y}_{2}\to\mathcal{Y}_{1}$ is a vector bundle of rank $r$, then
    $\mathcal{Y}_{2}\xrightarrow{\pi\circ\psi}\mathcal{X}]=[\mathcal{Y}_{1}\times_{k}\mathbb{A}^{r}_{k}\xrightarrow{\pi\circ p_{1}}\mathcal{X}].$
@@
We define the pullback and pushforward homomorphisms for stacks of finite type. Similarly to the content of [FSS18, Section 2.1], for a morphism $f\colon\mathcal{X}\to\mathcal{Y}$ of stacks of finite type possibly over different fields, we define the pullback homomorphism $f^{\ast}\colon\operatorname{Mot}(\mathcal{Y})\to\operatorname{Mot}(\mathcal{X})$ by
@@
The pullback homomorphism is well-defined because the surjective and universally injective morphisms are stable under base change, see [SP18, Tag03MH and Tag03MW].
For a morphism $f\colon\mathcal{X}\to\mathcal{Y}$ of finite type, we also define the pushforward homomorphism $f_{!}\colon\operatorname{Mot}(\mathcal{X})\to\operatorname{Mot}(\mathcal{Y})$ by
@@
The pullbacks and pushforwards satisfy the same properties as in [FSS18, Section 2.1]. For $k$-stacks $\mathcal{X}$ and $\mathcal{Y}$ of finite type, we have an external product $\boxtimes\colon\operatorname{Mot}(\mathcal{X})\otimes_{\mathbb{Z}}\operatorname{Mot}(\mathcal{Y})\to\operatorname{Mot}(\mathcal{X}\times_{k}\mathcal{Y})$.
@@
Let $\mathcal{X}$ be a stack locally of finite type. We define
@@
(2.1) $\operatorname{Mot}^{fin}(\mathcal{X})\coloneqq\cup_{\mathcal{U}\in\operatorname{Opf}(\mathcal{X})}(j_{\mathcal{U}})_{!}\left(\operatorname{Mot}(\mathcal{U})\right)\subset\operatorname{Mot}(\mathcal{X}),$
@@
where the limit is taken over the partially ordered set $\operatorname{Opf}(\mathcal{X})$. Note that the pullback $f^{\ast}$ and pushforward $f_{!}$ can be extended to the completions. The group $\overline{\operatorname{Mot}}^{fin}(\mathcal{X})$ is defined similarly to (2.1).